# Streamlined Supergravity

### v3

@@ 1 Introduction
Supergravity textbooks Wess:1992cp ; Freedman:2012zz and a more recent “Simple Supergravity” review DallAgata:2022nfr describe supergravities with linearly realized supersymmetry. The models with non-linearly realized supersymmetry have been actively studied over the last couple of decades, particularly in the context of cosmological applications of supergravity. These are best described as supergravity models that, in addition to unconstrained physical chiral multiplets $(Z^{i},\bar{Z}^{\bar{i}})$, contain one nilpotent chiral multiplet $(S,\bar{S})$ with $S^{2}=0$. This is a supergravity version of the anti-D3 brane uplifting of the potential in the KKLT model Kachru:2003aw required to produce de Sitter vacua in string theory. An early description of these models for inflation and dark energy can be found in Ferrara:2014kva , and in a more recent review Antoniadis:2024hvw .
@@
An important advantage of having the nilpotent multiplet in addition to standard chiral multiplets in the theory was emphasized in Kallosh:2015sea . The potential of these models at the vanishing scalar $s$ of the nilpotent multiplet has the form
@@
Important steps in developing supergravity with one nilpotent multiplet in addition to standard chiral multiplets where made over the years in Kallosh:2015tea ; Kallosh:2015pho ; Ferrara:2016een ; McDonough:2016der ; Freedman:2017obq ; Kallosh:2017wnt ; Achucarro:2017ing ; Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha ; Kallosh:2025jsb . A related new construction called “Liberated Supergravity” was developed in Farakos:2018sgq based on linearly realized off-shell supersymmetry and Kähler - Weyl invariance.
@@ 2 Construction of streamlined supergravity with a nilpotent superfield
Streamlining, by definition, means making something more efficient and simple by removing unnecessary parts. Note that the nilpotent sector acts behind the scenes; our only goal is to find a simple and efficient way to construct the most general scalar potentials and Kähler potentials in the physical sector. In our case, it involves two simplifications in the nilpotent sector of the theory.
@@
    A more general case would be the one where we add the $z,\bar{z}$-dependent terms that are linear in $s$ and $\bar{s}$. The result is known Schillo:2015ssx , supergravity is more complicated. But for our purposes, such complications are not warranted: we can reach our main goal without adding these extra terms in the nilpotent sector of the theory.
@@
In case of global supersymmetry we have $n$ chiral superfields $Z^{i}=z^{i}+\sqrt{2}\,\theta\psi^{i}_{z}+\theta^{2}f_{z}^{i}$ and one nilpotent superfield $S$, where $S=s+\sqrt{2}\,\theta\psi^{s}+\theta^{2}f_{s}$ and $S^{2}=0$. The nilpotency condition is resolved as follows for $f_{s}\neq 0$ Komargodski:2009rz
@@
This means that we have to impose a condition $s=0$ in the bosonic part of the action.
In the case of local supersymmetry, the resolution of the nilpotency condition is a generalization of the one in global supersymmetry. Superconformal theory with a nilpotent multiplet was studied in Kallosh:2015sea ; Kallosh:2015tea . Transition to supergravity followed the steps in Freedman:2012zz for usual chiral and vector multiplets, and special attention was given to the presence of the nilpotent chiral multiplet. As a result, the resolution of the nilpotency condition was established in Kallosh:2015sea ; Kallosh:2015tea and it takes the following form
@@
Here $\dots$ denote various terms containing fermions. We will present more details in the Appendix to describe the complete supergravity action, including scalars, fermions, and vectors. In supergravity in Kallosh:2015sea and here in eq. (2) in absence of fermions
@@
We will sometimes call it “new” rather than “streamlined” for simplicity. The resulting bosonic action of scalars is
@@ 3 Consistency issue of streamlined supergravity and its resolution
The potential at some point in the moduli space $T_{min}$ may have a de Sitter minimum where it is equal to $\Lambda\ll 1$ and we find that at this point
@@
This was viewed for a while in Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha as a restriction on the Kähler curvature of the moduli space, in this case, with $\mathcal{R}_{K}=-{2\over 3\alpha}$, the restriction was $|\mathcal{R}_{K}|<{2\over 3}$.
More recently, this issue was revisited in Kallosh:2025jsb , taking into account that $S$ is a nilpotent chiral multiplet. Consider the full supergravity action, which includes the Kähler metric $K_{s\bar{s}}$, which could lead to an inconsistency because it is not positive definite. These are kinetic terms in the action of the form
@@
These terms do not require $K_{s\bar{s}}$ to be positive. However, the second term in eq. (18) is quadratic in spinors, and its presence in the action indicates a possible inconsistency of the theory, as always in the case of the negative Kähler metric.
However, we can use local supersymmetry to gauge-fix the fermion from the nilpotent multiplet to vanish,
@@
This is a unitary gauge introduced in Bergshoeff:2015tra in supergravity interacting with one nilpotent multiplet and in Kallosh:2015sea in supergravity interacting with a nilpotent multiplet and other chiral multiplets. In this gauge, the second term in eq. (18) is absent, so the problem is solved, both terms in eq. (18) are absent from the gauge-fixed action.
@@
The requirement that goldstino $v$ is vanishing means in our new supergravity models that
@@
Therefore, the kinetic terms of the boson $s$ and the fermion $\psi^{s}$ with a negative $K_{s\bar{s}}$ are not harmful. There is no vacuum instability due to these fermions since their kinetic term is either absent in $\psi^{s}=0$ gauge, or involves terms with negative $K_{s\bar{s}}$ which are not quadratic in physical fermions since they contain coupling to moduli $z^{i}$. Thus the streamlined supergravity defined by $K_{s\bar{s}}$ in eqs. (11) is consistent even in the models where $K_{s\bar{s}}<0$ at some points in moduli space.
@@ Appendix A Complete streamlined supergravity with scalars, fermions, and vectors
This complete action with chiral and vector multiplets is complicated before gauge-fixing local supersymmetry. It is complicated before gauge-fixing, even in the case of pure de Sitter supergravity Bergshoeff:2015tra . But in the unitary gauge with $\psi^{s}=0$, both the action of the pure de Sitter supergravity in Bergshoeff:2015tra as well as the action with chiral and vector multiplets become very simple. In our case here, we find in the unitary gauge with $\psi^{s}=0$
@@
Finally, we would like to note here that particular choice of $K_{s\bar{s}}$ and $F_{s}$ in our scalar potential in (2) $e^{K}|F_{s}|^{2}K^{s\bar{s}}$ reproduces the potential (1.1) in Cribiori:2017laj . Namely, with $K_{s\bar{s}}=e^{K/3},|F_{s}|^{2}={\xi^{2}\over 2}$, we find that our potential is equal to ${\xi^{2}\over 2}e^{2K\over 3}$. This potential was interpreted in Cribiori:2017laj in the context of the gauge-fixed superconformal theory as a Fayet-Iliopoulos term in supergravity without gauged R-symmetry. It was, however, also observed there that in the context of constrained multiplets, this potential originates from the superpotential $W=W(z)-{\xi\over\sqrt{2}}s$. This is in agreement with our term F-term potential $e^{K}|F_{s}|^{2}K^{s\bar{s}}$ with $K_{s\bar{s}}=e^{K/3}$, $|F_{s}|^{2}={\xi^{2}\over 2}$.
@@ Appendix B $SL(2,\mathbb{Z})$ invariant supergravity
We consider supergravity action with a physical chiral superfield where its scalar $\tau=iT$ is a modulus of the upper half-plane, and there is a nilpotent superfield with a scalar $s$ and fermion $\psi^{s}$. We study the action, including fermions, gravitino, and the fermion partner of $T$, which we call $\psi^{T}$, in the unitary gauge $\psi^{s}=0$. According to eq. (45), this means we have to start with standard supergravity with 2 chiral superfields with scalars $T,s$, and take into account that the potential (46) differs from the standard supergravity potential with only one superfield $T$.
@@
Now, as different from Ferrara:1989bc , we start the action with two multiplets: the physical one and the nilpotent one. Gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})$ and its $T$ derivatives, where $G(T,\bar{T};s,\bar{s})=G(T,\bar{T})+\dots$ Here $\dots$ means terms linear in $s$ and $\bar{s}$ and in $s\bar{s}$. But according to eq. (45) in the unitary gauge $\psi^{s}=0$, we have to take the standard action for both multiplets at $\psi^{s}=0,s=0$. Thus, in the final action, gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})\Big|_{s=0}=G(T,\bar{T})$ and its derivatives in $T,\bar{T}$ as shown in eq. (7) in Ferrara:1989bc .

### v4

@@ 1 Introduction
Supergravity textbooks Wess:1992cp ; Freedman:2012zz and a more recent “Simple Supergravity” review DallAgata:2022nfr describe supergravities with linearly realized supersymmetry. The models with non-linearly realized supersymmetry have been actively studied over the last couple of decades, particularly in the context of cosmological applications of supergravity. These are best described as supergravity models that, in addition to unconstrained physical chiral multiplets $(Z^{i},\bar{Z}^{\bar{i}})$, contain one nilpotent chiral multiplet $(S,\bar{S})$ with $S^{2}=0$. This is a supergravity version of the anti-D3 brane uplifting of the potential in the KKLT model Kachru:2003aw , which is required to produce de Sitter vacua in string theory. Early descriptions of these models for inflation and dark energy can be found in Ferrara:2014kva and in a more recent review Antoniadis:2024hvw .
@@
An important advantage of including the nilpotent multiplet, in addition to standard chiral multiplets, in the theory was emphasized in Kallosh:2015sea . The potential of these models at the vanishing scalar $s$ of the nilpotent multiplet has the form
@@
Important steps in developing supergravity with one nilpotent multiplet in addition to standard chiral multiplets where made over the years, see for example Antoniadis:2014oya ; DallAgata:2014qsj ; Dudas:2015eha ; Kallosh:2015tea ; Kallosh:2015pho ; Ferrara:2016een ; McDonough:2016der ; Freedman:2017obq ; Kallosh:2017wnt ; Achucarro:2017ing ; Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha ; Kallosh:2025jsb . A related new construction called “Liberated Supergravity” was developed in Farakos:2018sgq based on linearly realized off-shell supersymmetry and Kähler - Weyl invariance.
@@ 2 Construction of streamlined supergravity with a nilpotent superfield
Streamlining, by definition, means making something more efficient and simpler by removing unnecessary parts. Note that the nilpotent sector acts behind the scenes; our only goal is to find a simple and efficient way to construct the most general scalar potentials and Kähler potentials in the physical sector. In our case, it involves two simplifications in the nilpotent sector of the theory.
@@
    A more general case would be the one where we add the $z,\bar{z}$-dependent terms that are linear in $s$ and $\bar{s}$. The result is known Schillo:2015ssx , supergravity is more complicated. However, for our purposes, such complications are not warranted: we can achieve our main goal without introducing these additional terms in the nilpotent sector of the theory.
@@
The original Volkov-Akulov theory of nonlinearly realized global supersymmetry Volkov:1972jx did not rely on models with linearly realized supersymmetry with unconstrained chiral multiplets, since the Volkov-Akulov model was discovered earlier. In case of global supersymmetry we have $n$ chiral superfields $Z^{i}=z^{i}+\sqrt{2}\,\theta\psi^{i}_{z}+\theta^{2}f_{z}^{i}$ and one nilpotent superfield $S$, where $S=s+\sqrt{2}\,\theta\psi^{s}+\theta^{2}f_{s}$ and $S^{2}=0$. The nilpotency condition is resolved as follows, for $f_{s}\neq 0$ Casalbuoni:1988xh ; Komargodski:2009rz ,
@@
which brings the original theory with linear supersymmetry to a Volkov-Akulov type model with a non-linearly realized supersymmetry. It was observed in Casalbuoni:1988xh that the nilpotency condition $S^{2}=0$ is achieved in the model with linear supersymmetry in the limit of infinite scalar mass. This limit is performed by sending the vacuum curvature of the Kähler manifold to infinity. The nilpotency constraint (7) means that we have to impose a condition $s=0$ in the bosonic part of the action.
In the case of local supersymmetry, the resolution of the nilpotency condition is a generalization of the one in global supersymmetry. Superconformal theory with a nilpotent multiplet was studied in Kallosh:2015sea ; Kallosh:2015tea . Transition to supergravity followed the steps described in Freedman:2012zz for usual chiral and vector multiplets, and special attention was given to the presence of the nilpotent chiral multiplet. As a result, the resolution of the nilpotency condition was established in Kallosh:2015sea ; Kallosh:2015tea . It takes the following form:
@@
Here, $\dots$ denotes various terms containing fermions. We will provide further details in the Appendix, describing the complete supergravity action, including scalars, fermions, and vectors. In supergravity in Kallosh:2015sea and here in eq. (2) in absence of fermions
@@
We will sometimes refer to it as “new” rather than “streamlined” for simplicity. The resulting bosonic action of scalars is
@@ 3 Consistency issue of streamlined supergravity and its resolution
The potential at some point in the moduli space $T_{min}$ may have a de Sitter minimum where it is equal to $\Lambda\ll 1$, and we find that at this point
@@
This was viewed for a while in Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha as a restriction on the Kähler curvature of the moduli space, in this case, with $\mathcal{R}_{K}=-{2\over 3\alpha}$, the restriction was $|\mathcal{R}_{K}|<{2\over 3}$. More recently, this issue was revisited in Kallosh:2025jsb , taking into account that $S$ is a nilpotent chiral multiplet. We will describe this case in a more general setting below. But first, we would like to discuss different points of view on the origin of non-linearly realized supersymmetry.
There was a trend, following Casalbuoni:1988xh ; Komargodski:2009rz , to obtain the nilpotent constraint in global supersymmetry models by starting with a linearly realized supersymmetry, giving a large mass to the scalar field, and integrating it out. Analogous ideas were also developed in the context of supergravity. This procedure may be helpful for the reinterpretation of certain aspects of nonlinearly realized global supersymmetry. In particular, starting with linear supersymmetry models, one can evaluate the path integral in the limit of infinite sgoldstino mass Kallosh:2015pho to derive de Sitter supergravity Bergshoeff:2015tra ; Hasegawa:2015bza , or investigate the decoupling limits to derive a model with a nilpotent multiplet; see, for example, Farakos:2013ih ; DallAgata:2016syy . A consistency of this limiting procedure relies on the consistency of the theory with the linearly realized supersymmetry model before the formal limit to a nilpotent multiplet with nonlinear supersymmetry is taken. This means, in particular, that one can establish a relation between supergravity with nonlinearly realized supersymmetry and with linearly realized supersymmetry only for the models with $K_{s\bar{s}}>0$.
However, we would like to remind here that supergravity theory with a nilpotent multiplet was derived from the underlying superconformal theory in Bergshoeff:2015tra ; Hasegawa:2015bza ; Kallosh:2015sea ; Ferrara:2016een . This theory was not derived as a decoupling limit of supergravity with linearly realized supersymmetry, and its internal consistency does not rely on any relation between the linear and nonlinear realizations of supersymmetry in the large-mass limit.
Moreover, the existence of the D-branes shows that a non-linearly realized supersymmetry does not have to be a limit of the kind considered in Casalbuoni:1988xh ; Komargodski:2009rz . A supersymmetric Born-Infeld theory describing a D9-brane has a geometric form of the Volkov-Akulov-type Kallosh:1997aw . There are Volkov-Akulov fermions living on the $\overline{D3}$ brane Bergshoeff:2015jxa , which is a part of the KKLT mechanism of uplifting AdS vacua to dS vacua. The uplifting $\overline{D3}$ brane is either present or not, which means that non-linearly realized supersymmetry is either present or not; it was not constructed by a limiting procedure. A recent review of these topics can be found in Dudas:2025ubq .
We will study below the issue of consistency of the supergravity with a nilpotent multiplet Bergshoeff:2015tra ; Hasegawa:2015bza ; Kallosh:2015sea ; Ferrara:2016een without attempting to re-interpret it in terms of the theory with the linear supersymmetry, where the decoupling limit is required. We will find out that the sign of the Kähler metric of the nilpotent multiplet does not matter: bosons and fermions of the nilpotent multiplet are not excited in the unitary gauge. In this sense, the situation here is very similar to that in other gauge theories, where ghosts and tachyons that appear at intermediate stages of calculations are harmless if they disappear in the unitary gauge.
To investigate this issue in detail, we consider the full supergravity action, which includes the Kähler metric $K_{s\bar{s}}$, which could lead to an inconsistency because it is not positive definite. These are kinetic terms in the action of the form
@@
According to Bellazzini:2016xrt , this term is subject to positivity constraints, which would be inconsistent with the negative metric. Eqs. (5.3), (5.5) in Bellazzini:2016xrt show that in the case of global supersymmetry studied in Casalbuoni:1988xh ; Komargodski:2009rz ; Bellazzini:2016xrt indeed, one has to require that $K_{s\bar{s}}>0$. The second term in eq. (18) is quadratic in spinors, and its presence in the action directly indicates a possible inconsistency of the theory, as always in the case of the negative Kähler metric in quadratic kinetic terms.
However, in the case of a supergravity with a nilpotent multiplet (but not in the case of a global supersymmetry), we can use local supersymmetry to gauge-fix the fermion from the nilpotent multiplet to vanish,
@@
This is the unitary gauge, which was introduced in Bergshoeff:2015tra in supergravity interacting with one nilpotent multiplet and in Kallosh:2015sea in supergravity interacting with a nilpotent multiplet and other chiral multiplets. In this gauge, both terms in eq. (18), with account of eq. (19), are absent from the gauge-fixed action, so the problem is solved. A negative Kähler metric for the nilpotent multiplet does not cause a physical instability of the system. This eliminates the corresponding constraints on the parameters of the cosmological theories of the kind we described in the example in eqs. (14)-(17).
@@
The requirement that goldstino $v$ vanishes in our new supergravity models means that
@@
Therefore, the kinetic terms of the boson $s$ and the fermion $\psi^{s}$ with a negative $K_{s\bar{s}}$ are not harmful. There is no vacuum instability due to $s$ and $\psi^{s}$ since their kinetic terms are either absent in $\psi^{s}=0$ gauge, or in $v=0$ gauge involve terms with negative $K_{s\bar{s}}$ which are not quadratic in physical fermions $\psi^{i}_{z}$ since they contain coupling to moduli $z^{i}$. Thus, the streamlined supergravity defined by $K_{s\bar{s}}$ in eqs. (11) is consistent even in the models where $K_{s\bar{s}}<0$ at some points in moduli space.
@@ Appendix A Complete streamlined supergravity with scalars, fermions, and vectors
This complete action, including chiral and vector multiplets, is complicated by local supersymmetry before gauge fixing. It is complicated even in the case of pure de Sitter supergravity Bergshoeff:2015tra . But in the unitary gauge with $\psi^{s}=0$, both the action of the pure de Sitter supergravity in Bergshoeff:2015tra as well as the action with chiral and vector multiplets become very simple. In our case here, we find in the unitary gauge with $\psi^{s}=0$
@@
Finally, we would like to note here that particular choice of $K_{s\bar{s}}$ and $F_{s}$ in our scalar potential in (2) $e^{K}|F_{s}|^{2}K^{s\bar{s}}$ reproduces the potential (1.1) in Cribiori:2017laj . Namely, with $K_{s\bar{s}}=e^{K/3},|F_{s}|^{2}={\xi^{2}\over 2}$, we find that our potential is equal to ${\xi^{2}\over 2}e^{2K\over 3}$. This potential was interpreted in Cribiori:2017laj in the context of the gauge-fixed superconformal theory as a Fayet-Iliopoulos term in supergravity without gauged R-symmetry. It was, however, also observed there that ,in the context of constrained multiplets, this potential originates from the superpotential $W=W(z)-{\xi\over\sqrt{2}}s$. This is in agreement with our term F-term potential $e^{K}|F_{s}|^{2}K^{s\bar{s}}$ with $K_{s\bar{s}}=e^{K/3}$, $|F_{s}|^{2}={\xi^{2}\over 2}$.
@@ Appendix B $SL(2,\mathbb{Z})$ invariant supergravity
We consider supergravity action with a physical chiral superfield where its scalar $\tau=iT$ is a modulus of the upper half-plane, and there is a nilpotent superfield with a scalar $s$ and fermion $\psi^{s}$. We study the action, including fermions, gravitino, and the fermion partner of $T$, which we call $\psi^{T}$, in the unitary gauge $\psi^{s}=0$. According to eq. (45), this means that we have to start with standard supergravity with 2 chiral superfields with scalars $T,s$, and take into account that the potential (46) differs from the standard supergravity potential with only one superfield $T$.
@@
Now, as different from Ferrara:1989bc , we begin with an action containing two multiplets: the physical multiplet and the nilpotent multiplet. Gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})$ and its $T$ derivatives, where $G(T,\bar{T};s,\bar{s})=G(T,\bar{T})+\dots$ Here $\dots$ means terms linear in $s$ and $\bar{s}$ and in $s\bar{s}$. But according to eq. (45) in the unitary gauge $\psi^{s}=0$, we have to take the standard action for both multiplets at $\psi^{s}=0,s=0$. Thus, in the final action, gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})\Big|_{s=0}=G(T,\bar{T})$ and its derivatives in $T,\bar{T}$ as shown in eq. (7) in Ferrara:1989bc .